Consider the function g defined by:


g(x,y) = cos ( "\u221ay" ) + 1/log3(x - y)


Determine ∂2g/∂y∂x



1
Expert's answer
2022-05-10T23:11:10-0400

Function "g(x,y)" has the form: "g(x,y)=\cos(x\pi\sqrt{y})+\frac{1}{log_3(x-y)}=\cos(x\pi\sqrt{y})+\frac{\ln(3)}{ln(x-y)}".

We receive:

"\frac{\partial g}{\partial y}=-\frac{x\pi}{2\sqrt{y}}\sin(x\pi\sqrt{y})+\frac{\ln(3)}{(\ln(x-y))^2}\cdot\frac{1}{x-y}\\". "\,\frac{\partial^2g}{\partial y\partial x}=-\frac{\pi}{2\sqrt{y}}\sin(x\pi\sqrt{y})-\frac{x\pi^2}{2}\cos(x\pi\sqrt{y})-\frac{(\ln(3))(2+\ln(x-y))}{(\ln(x-y))^3}\cdot\frac{1}{(x-y)^2}".

Answer: "\frac{\partial^2g}{\partial y\partial x}=-\frac{\pi}{2\sqrt{y}}\sin(x\pi\sqrt{y})-\frac{x\pi^2}{2}\cos(x\pi\sqrt{y})-\frac{(\ln(3))(2+\ln(x-y))}{(\ln(x-y))^3}\cdot\frac{1}{(x-y)^2}"